# Large deviations for the smallest eigenvalue of a deformed GOE with an outlier

### v1

@@ 2. Large deviations in the pure point case
Let $B_{N}$ be a $N\times N$ diagonal matrix with eigenvalues $\Lambda$ with multiplicity $1$ and $\eta_{1},\ldots,\eta_{p}$ with multiplicity $N_{1},\cdots,N_{p}$ such that that for each $i\in\{1,\ldots,p\}$, $N_{i}/N$ converges towards $\alpha_{i}>0$. Let $G_{N}$ be a GOE matrix of size $N\times N$ and variance $t>0$. We assume that $\Lambda<\eta_{1}<\eta_{2}<\cdots<\eta_{p}$ and denote
@@
Let $\Delta^{p+1}$ be the $(p+1)$-simplex as defined in (1.14). We recall that $\nu\boxplus\sigma_{t}$ is the limiting distribution of the spectrum of $G_{N}+B_{N}$ and $\ell_{\nu,t}$ denotes the leftmost point in the support of $\nu\boxplus\sigma_{t}$. The goal of this section is to prove the following large deviation principle, which specializes Theorem 1.2 to the case where $B_{N}$ has a fixed number of distinct eigenvalues. Let $I_{\nu,t}^{\Lambda}$ be as defined in (1.9).
@@
It is easy to see that $I_{\nu,t}^{\Lambda}$ is a good rate function. Indeed, it is continuous on $(-\infty,\ell_{\nu,t})$ since $S_{\nu}$ is continuous on $(-\infty,\ell_{\nu})$, $\omega_{\nu,t}$ is continuous and increasing on $(-\infty,\ell_{\nu,t})$, with values in $(-\infty,\omega_{\nu,t}(\ell_{\nu,t})]$, and $\omega_{\nu,t}^{*}$ is continuous with values in $[\omega_{\nu,t}(\ell_{\nu}),\ell_{\nu,t}]$. This shows that $I_{\nu,t}^{\Lambda}$ is continuous expect possibly at $\ell_{\nu,t}$ where it is lower semicontinuous. Moreover, $I_{\nu,t}^{\Lambda}(\ell^{\Lambda}_{\nu,t})=0$. Indeed, if $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$, then $\ell_{\nu,t}^{\Lambda}=\rho_{\nu,t}^{\Lambda}$ and $\omega_{\nu,t}(\rho^{\Lambda}_{\nu,t})=\Lambda=\gamma(\lambda,\Lambda)$. Besides, if $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$, $\ell_{\nu,t}^{\Lambda}=\ell_{\nu,t}$ and $\omega_{\nu,t}(\ell_{\nu,t})=\omega_{\nu,t}^{*}(\ell_{\nu,t})=\gamma(\lambda,\Lambda)$. Let us show that
$\lim_{\lambda\to-\infty}{I_{\nu,t}^{\Lambda}}(\lambda)=+\infty.$
As $\lambda\to-\infty$, we can notice that $\omega_{\nu,t}(\lambda)\sim\lambda$. It follows that
@@
Moreover, the fact that $\lambda_{1}$ is exponentially tight is clear since $B_{N}$ is uniformly bounded and the smallest eigenvalue of the GOE is well known to be exponentially tight [AGZ10, Section 2.6.2]. We more precisely find, see the proof of [GH20, Lemma 1.8], that there exists a finite constant $C$ such that for $N$ large enough,
@@
The main point of the proof, on which we will focus for the remainder of this section, is therefore to prove a weak large deviation principle, namely that for every $\lambda<\ell_{\nu,t}$,
@@
Indeed the large deviation principle (2.3) is easy to prove for $x>\ell_{\nu,t}$. Indeed, by concentration of measure, we know that for every $\zeta>0$, there exists $C_{\zeta}>0$ such that for $N$ large enough,
@@
which corresponds to the announced statement, since by Lemma B.1, $I_{\nu,t}^{\Lambda}(\ell_{\nu,t}^{\Lambda})=0$.
If ${\ell_{\nu,t}}>\ell^{\Lambda}_{\nu,t}$, the monotonicity of the large deviation rate function proven in Lemma B.1 implies that for every $\delta>0$ small enough, every $\gamma<\delta/4$,
$\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\ell_{\nu,t}|\leq\delta)\leq\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\ell_{\nu,t}-\delta|\leq\gamma)$
so that (2.3) at $\lambda=\ell_{\nu,t}-\delta$ gives, after taking the limit $\gamma\rightarrow 0$,
$\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\ell_{\nu,t}|\leq\delta)\leq-I_{\nu,t}^{\Lambda}(\ell_{\nu,t}-\delta)$
and the left-continuity of $I_{\nu,t}^{\Lambda}$ at $\ell_{\nu,t}$ implies
$\lim_{\delta\rightarrow 0}\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\ell_{\nu,t}|\leq\delta)\leq-I_{\nu,t}^{\Lambda}(\ell_{\nu,t})\,.$ (2.5)
@@
which implies by (2.3) that
@@ 2.1. Projection onto the first eigenvector
where $p_{v_{1}}^{\perp}$ is the project onto the orthocomplement of $v_{1}$.
@@
Let $w_{0}$ be a unitary eigenvector of $B_{N}$ associated to $\Lambda$. For each $k=1,\ldots,p$, let us denote by $P_{E_{\eta_{k}}}$ the orthogonal projection onto the eigenspace $E_{\eta_{k}}$ corresponding to the eigenvalue $\eta_{k}$. The quantities of Lemma 2.2 can then be written as follows:
$\langle v_{1},B_{N}v_{1}\rangle$ $=\Lambda\langle v_{1},w_{0}\rangle^{2}+\sum_{k=1}^{p}\eta_{k}\|P_{E_{\eta_{k}}}v_{1}\|^{2}$ (2.8)
$\langle v_{1},B_{N}^{2}v_{1}\rangle$ $=\Lambda^{2}\langle v_{1},w_{0}\rangle^{2}+\sum_{k=1}^{p}\eta_{k}^{2}\|P_{E_{\eta_{k}}}v_{1}\|^{2}.$ (2.9)
Define $Y^{N}=(Y_{0}^{N},\ldots,Y_{k}^{N})$ by
The variable $Y^{N}$ takes values into the simplex $\Delta^{p+1}$.
@@ 2.2. LDP for Dirichlet variables
One can observe that the random variable $Y^{N}$ follows a Dirichlet law with parameter $(1,N_{1},\ldots,N_{p})$ which are such that $N_{i}/N$ goes to $\alpha_{i}>0$ for each $i=1,\ldots,p$. Using the explicit density of the Dirichlet distribution, it easy to prove [GH22] by Laplace’s method that $Y^{N}$ satisfies the following LDP:
@@ 2.3. Outliers of $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$
As outlined in Subsection 1.4, the proof involves computing the probability that the smallest eigenvalue of $G_{N-1}+p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ is larger than $x$, where $G_{N-1}$ is a GOE matrix of size $(N-1)\times(N-1)$ on the orthocomplement of $v_{1}$, for given $x\leq\ell_{\nu,t}$ and $Y\in\Delta^{p+1}$. Thus, we first study the outliers of the matrix $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ when $v_{1}$ is such that $Y_{N}$ is equal to some $Y$, and then invoke known results on the BBP transition.
@@
Let $B_{N}$ be a $N\times N$ diagonal matrix with eigenvalues $\eta_{0}:=\Lambda<\eta_{1}<\cdots<\eta_{p}$ with multiplicity $N_{0}=1,N_{1},\cdots,N_{p}$ such that that for each $i\in\{1,\ldots,p\}$, $N_{i}\geq 2$. For $Y\in\Delta^{p+1}$ so that $y_{0}\in(0,1)$, we set $\Phi(\Lambda,Y)\in[\Lambda,\eta_{1}]$ to be the smallest solution to the equation
When $y_{0}=0$, we set $\Phi(\Lambda,Y)=\Lambda$ and when $y_{0}=1$ we set $\Phi(\Lambda,Y)=\eta_{1}$. Then $\Phi(\Lambda,.)$ is continuous. Moreover:
@@
    Let $v_{1}\in\mathbb{R}^{N}$ be a unit vector, $v_{1}\notin\cup_{k=0}^{p}E_{\eta_{k}}$. For each $k=0,\ldots,p$, denote
    $y_{k}:=\|v_{1}\|_{E_{\eta_{k}}}^{2}.$ (2.13)
    Then the matrix $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ has eigenvalues $\Lambda_{1}<\cdots<\Lambda_{p}$ of multiplicities $1$ and $\eta_{1},\ldots\eta_{p}$ of multiplicities $N_{1}-1,\ldots,N_{p}-1$. Moreover for each $k=1,\ldots,p$, $\Lambda_{k}\in(\eta_{k-1},\eta_{k})$. Finally, $\Lambda_{1}=\Phi(\Lambda,Y)$.
@@
    Let $v_{1}$ be a unit vector such that $v_{1}\in E_{\eta_{k}}$ for some $k=0,\ldots,p$. Then the matrix $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ has eigenvalues $\eta_{i}$ with multiplicity $N_{i}$ for each $i\in\{0,\ldots,p\}\setminus\{k\}$, and $\eta_{k}$ with multiplicity $N_{k}-1$. Moreover, the smallest eigenvalue $\Lambda_{1}$ of $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ equals $\Phi(\Lambda,Y)$.
@@
Let us prove (1). Let $v_{1}$ be a unit vector, which is not an eigenvector of $B_{N}$. Let $B_{N}^{\prime}:=p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$. Let us show that $\eta_{k}$ has multiplicity $N_{k}-1$. The eigenspace of $B_{N}^{\prime}$ associated to $\eta_{k}$ is
@@
Since $v_{1}\notin E_{\eta_{k}}$, $E_{\eta_{k}}^{\prime}$ has dimension $N_{k}-1$. Moreover $\Lambda$ is not an eigenvalue of $B_{N}^{\prime}$. There are therefore $p$ new eigenvalues $\Lambda_{1}\leq\cdots\leq\Lambda_{p}$. Recall that if $\lambda_{1}\leq\cdots\leq\lambda_{N}$ are the eigenvalues of $B_{N}$ and $\lambda_{1}^{\prime}\leq\cdots\leq\lambda_{N}^{\prime}$ are the eigenvalues of $B_{N}^{\prime}$, then by the Courant-Fischer theorem, $\lambda$ and $\lambda^{\prime}$ are interlaced:
@@
Therefore for each $i=1,\ldots,p$, we have $\Lambda_{i}\in(\eta_{i-1},\eta_{i}).$
Moreover, by the Courant-Fischer theorem, $\Lambda_{1}$ can be expressed as
@@
Let $v\in\mathbb{R}^{N}$ be a minimizer in the above problem. By variational calculus, there exist $C_{1},C\in\mathbb{R}$ such that
$B_{N}v=Cv+C_{1}v_{1}.$
Since $\langle v,v_{1}\rangle=0$, we have
$\langle v,B_{N}v\rangle=C=\Lambda_{1}.$
Since $\Lambda_{1}\notin\{\lambda_{1},\ldots,\lambda_{N}\}$, we deduce that for each $i=1,\ldots,N$,
$v_{i}=C_{1}\frac{(v_{1})_{i}}{\lambda_{i}-\Lambda_{1}}.$
Thus $\Lambda_{1}$ satisfies
$\langle v,v_{1}\rangle=\sum_{i=1}^{N}\frac{(v_{1})_{i}^{2}}{\lambda_{i}-\Lambda_{1}}=0,$

### v2

@@ 2. Large deviations in the pure point case
Let $B_{N}$ be a $N\times N$ diagonal matrix with eigenvalues $\Lambda$ with multiplicity $1$ and $\eta_{1},\ldots,\eta_{p}$ with multiplicity $N_{1},\cdots,N_{p}$ such that that for every $i\in\{1,\ldots,p\}$, $N_{i}/N$ converges towards $\alpha_{i}>0$. Let $G_{N}$ be a GOE matrix of size $N\times N$ and variance $t>0$. We assume that $\Lambda<\eta_{1}<\eta_{2}<\cdots<\eta_{p}$ and denote
@@
Throughout this section we denote by $\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{N}$ the eigenvalues of $X_{N}=B_{N}+G_{N}$. We denote in short $\lambda^{N}=(\lambda_{1},\ldots,\lambda_{N})$.
Let $\Delta^{p+1}$ be the $(p+1)$-simplex as defined in (1.17). We recall that $\nu\boxplus\sigma_{t}$ is the limiting distribution of the spectrum of $G_{N}+B_{N}$, namely the almost sure limit of $\frac{1}{N}\sum_{i=1}^{N}\delta_{\lambda_{i}}$, and $\ell_{\nu,t}$ denotes the leftmost point in the support of $\nu\boxplus\sigma_{t}$. The goal of this section is to prove the following large deviation principle, which specializes Theorem 1.2 to the case where $B_{N}$ has a fixed number of distinct eigenvalues. Let $I_{\nu,t}^{\Lambda}$ be as defined in (1.9).
@@
It is easy to see that $I_{\nu,t}^{\Lambda}$ is a good rate function. Indeed, it is continuous on $(-\infty,\ell_{\nu,t})$ since $S_{\nu}$ is continuous on $(-\infty,\ell_{\nu})$, $\omega_{\nu,t}$ is continuous and increasing on $(-\infty,\ell_{\nu,t})$, with values in $(-\infty,\omega_{\nu,t}(\ell_{\nu,t})]$, and $\omega_{\nu,t}^{*}$ is continuous with values in $[\omega_{\nu,t}(\ell_{\nu}),\ell_{\nu,t}]$. This shows that $I_{\nu,t}^{\Lambda}$ is continuous expect possibly at $\ell_{\nu,t}$ where it is lower semicontinuous. Moreover, $I_{\nu,t}^{\Lambda}(\ell^{\Lambda}_{\nu,t})=0$. Indeed, if $\Lambda\leq\omega_{\nu,t}(\ell_{\nu,t})$, then $\ell_{\nu,t}^{\Lambda}=\rho_{\nu,t}^{\Lambda}$ and $\omega_{\nu,t}(\rho^{\Lambda}_{\nu,t})=\Lambda=\gamma(\lambda,\Lambda)$. Besides, if $\Lambda\geq\omega_{\nu,t}(\ell_{\nu,t})$, $\ell_{\nu,t}^{\Lambda}=\ell_{\nu,t}$ and $\omega_{\nu,t}(\ell_{\nu,t})=\omega_{\nu,t}^{*}(\ell_{\nu,t})=\gamma(\lambda,\Lambda)$. Let us show that its level sets are compact. As $\lambda\to-\infty$, we notice that $\omega_{\nu,t}(\lambda)\sim\lambda$. It follows that
@@
The fact that $\lambda_{1}$ is exponentially tight is clear since $B_{N}$ is uniformly bounded and the smallest eigenvalue of the GOE is well known to be exponentially tight [AGZ10, Section 2.6.2]. We more precisely find, see the proof of [GH20, Lemma 1.8], that there exists a finite constant $C$ such that for $N$ large enough,
@@
The main point of the proof of Proposition 2.1, on which we will focus for the remainder of this section, is therefore to prove a weak large deviation principle, namely that for every $\lambda<\ell_{\nu,t}$,
@@
In fact, the large deviation principle (2.3) is easy to prove for $x>\ell_{\nu,t}$. Indeed, by concentration of measure, we know that for every $\zeta>0$, there exists $C_{\zeta}>0$ such that for $N$ large enough,
@@
which corresponds to the announced statement, since by Proposition 2.1, $I_{\nu,t}^{\Lambda}(\ell_{\nu,t}^{\Lambda})=0$.
If ${\ell_{\nu,t}}>\ell^{\Lambda}_{\nu,t}$, the monotonicity of the large deviation rate function proven in Lemma B.1 implies that for every $\delta>0$ small enough, every $y<\ell_{\nu,t}$ so that $|y-\ell_{\nu,t}|<|y-\ell_{\nu,t}^{\Lambda}|/2$,
$\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\ell_{\nu,t}|\leq\delta)\leq\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-y|\leq\delta)$
so that (2.3) at $y<\ell_{\nu,t}$ gives, after taking the limit $\delta\rightarrow 0$,
$\limsup_{\delta\rightarrow 0}\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\ell_{\nu,t}|\leq\delta)\leq-I_{\nu,t}^{\Lambda}(y)$
and the left-continuity of $I_{\nu,t}^{\Lambda}$ at $\ell_{\nu,t}$ implies by letting $y$ going to $\ell_{\nu,t}$ that
$\limsup_{N\rightarrow\infty}\frac{1}{N}\log\mathbb{P}(|\lambda_{1}-\ell_{\nu,t}|\leq\delta)\leq-I_{\nu,t}^{\Lambda}(\ell_{\nu,t})\,.$ (2.5)
@@
which implies by (2.3), after letting $N$ going to infinity and $\gamma$ to zero, that
@@ 2.1. Projection onto the first eigenvector
where $p_{v_{1}}^{\perp}$ is the projection onto the orthocomplement of $v_{1}$. Moreover, let $w_{0}$ denotes a unitary eigenvector of $B_{N}$ associated to $\Lambda$ and $P_{E_{\eta_{k}}}$ the orthogonal projection onto the eigenspace $E_{\eta_{k}}$ corresponding to the eigenvalue $\eta_{k}$. Denote by $Y^{N}=(Y_{0}^{N},\ldots,Y_{p}^{N})$ the vector in $\Delta^{p+1}$ given by
Then,
$-2\lambda_{1}\langle v_{1},Mv_{1}\rangle+2\langle v_{1},M^{2}v_{1}\rangle-\langle v_{1},Mv_{1}\rangle^{2}=4tL^{\Lambda}_{\nu,t}(\lambda_{1},Y^{N})$ (2.8)
with $L_{\nu,t}^{\Lambda}:(-\infty,\ell_{\nu,t}]\times\Delta^{p+1}\to\mathbb{R}$, given for all $\lambda\in(-\infty,\ell_{\nu,t}]$ and $Y\in\Delta^{p+1}$ by
@@
Equality (2.6) follows. To check (2.8), observe that
$\langle v_{1},B_{N}v_{1}\rangle$ $=\Lambda\langle v_{1},w_{0}\rangle^{2}+\sum_{k=1}^{p}\eta_{k}\|P_{E_{\eta_{k}}}v_{1}\|^{2}=\Lambda Y_{0}^{N}+\sum_{k=1}^{p}\eta_{k}Y_{k}^{N},$ (2.11)
$\langle v_{1},B_{N}^{2}v_{1}\rangle$ $=\Lambda^{2}\langle v_{1},w_{0}\rangle^{2}+\sum_{k=1}^{p}\eta_{k}^{2}\|P_{E_{\eta_{k}}}v_{1}\|^{2}=\Lambda^{2}Y_{0}^{N}+\sum_{k=1}^{p}\eta_{k}^{2}Y_{k}^{N}.$ (2.12)
Equality (2.8) follows by plugging (2.12) in (2.6).
@@ 2.2. LDP for Dirichlet variables
One can observe that the random variable $Y^{N}$ defined in (2.7) follows a Dirichlet law with parameter $(1,N_{1},\ldots,N_{p})$ which are such that $N_{i}/N$ goes to $\alpha_{i}>0$ for every $i=1,\ldots,p$. Using the explicit density of the Dirichlet distribution, it is easy to prove by Laplace’s method [GH22] that $Y^{N}$ satisfies the following LDP:
@@ 2.3. Outliers of $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$
As outlined in Subsection 1.4, the proof involves computing the probability that the smallest eigenvalue of $G_{N-1}+B_{N-1}^{\prime}$ is larger than $x$, where $G_{N-1}$ is a GOE matrix of size $(N-1)\times(N-1)$ on the orthocomplement of $v_{1}$ and $B_{N-1}^{\prime}$ is the $(N-1)\times(N-1)$ matrix given by $p_{v_{1}}^{\perp}B_{N}p_{v_{1}}^{\perp}$ in $v_{1}^{\perp}$. Thus, we first study the outliers of the matrix $B_{N-1}^{\prime}$ when $v_{1}$ is such that $Y_{N}$ is equal to some $Y$, and then invoke known results on the BBP transition. Here, $x\leq\ell_{\nu,t}$ and $Y\in\Delta^{p+1}$ are given.
We define $\Phi(\Lambda,\cdot):\Delta^{p+1}\to\mathbb{R}$ by setting for all $Y\in\Delta^{p+1}$,
$\Phi(\Lambda,Y)=\begin{cases}\Lambda&\text{if $y_{0}=0$}\\ \Lambda_{1}&\text{if $y_{0}\in(0,1],y_{1}\in(0,1]$}\\ \eta_{1}&\text{if $y_{1}=0$},\end{cases}$ (2.14)
where $\Lambda_{1}\in[\Lambda,\eta_{1}]$ is the unique solution in $[\Lambda,\eta_{1}]$ of the equation
Observe that $\Lambda_{1}$ is always uniquely defined when $y_{0}$ and $y_{1}$ are positive since (2.15) is equivalent to
$f(\Lambda_{1})=1\quad\mbox{ with }\quad f(x)=\sum_{k=1}^{p}y_{k}\frac{\eta_{k}-\Lambda}{\eta_{k}-x}\,.$ (2.16)
Clearly $f$ is strictly increasing on $[\Lambda,\eta_{1}]$ and $f([\Lambda,\eta_{1}])=[\sum_{k=1}^{p}y_{k},(+\infty)1_{y_{1}>0}+(\sum_{k=2}^{p}y_{k}\frac{\eta_{k}-\Lambda}{\eta_{k}-\eta_{1}})1_{y_{1}=0}]$. Hence, since $\sum_{k=1}^{p}y_{k}=1-y_{0}<1$ if $y_{0}>0$, we see that (2.16) has a unique solution when $y_{0}$ and $y_{1}$ are positive. Note also that $y_{0}=1$ implies $y_{1}=0$ and therefore $\Phi(\Lambda,Y)=\eta_{1}$. Moreover, when $y_{1}=0$, there might be no solution if $\sum_{k=2}^{p}y_{k}\frac{\eta_{k}-\Lambda}{\eta_{k}-\eta_{1}}<1$, for instance if $y_{0}$ is big enough and $y_{1}$ equal to zero.
@@
Let $B_{N}$ be a $N\times N$ diagonal matrix with eigenvalues $\eta_{0}:=\Lambda<\eta_{1}<\cdots<\eta_{p}$ with multiplicity $N_{0}=1,N_{1},\cdots,N_{p}$ such that that for every $i\in\{1,\ldots,p\}$, $N_{i}\geq 2$. Let $v_{1}\in\mathbb{R}^{N}$ be a unit vector and $B_{N-1}^{\prime}$ be the $(N-1)\times(N-1)$ matrix given by the restriction of $B_{N}$ to the orthocomplement of $v_{1}$. Moreover:
@@
    Let $\kappa>0$. Let $\Phi(\Lambda,Y)$ be the unique solution of (2.14) for $Y\in[0,1]\times[\kappa,1]\times[0,1]^{p-1}\cap\Delta^{p+1}$. Then $\Phi(\Lambda,.)$ is continuous on $[0,1]\times[\kappa,1]\times[0,1]^{p-1}\cap\Delta^{p+1}$.
@@
    Assume $v_{1}\notin\cup_{k=0}^{p}E_{\eta_{k}}$. For each $k=0,\ldots,p$, denote
    $y_{k}(v_{1}):=\|v_{1}\|_{E_{\eta_{k}}}^{2}.$ (2.17)
    Let $Y(v_{1}):=(y_{0}(v_{1}),\ldots,y_{p}(v_{1}))$ and assume $y_{1}(v_{1})>0$. Then, the matrix $B_{N-1}^{\prime}$ has eigenvalues $\Lambda_{1}<\cdots<\Lambda_{p}$ of multiplicities $1$ and $\eta_{1},\ldots\eta_{p}$ of multiplicities $N_{1}-1,\ldots,N_{p}-1$. Moreover for every $k=1,\ldots,p$, $\Lambda_{k}\in(\eta_{k-1},\eta_{k})$. Finally, $\Lambda_{1}=\Phi(\Lambda,Y(v_{1}))$.
  3. (3)
    If $v_{1}\in E_{\eta_{k}}$ for some $k=0,\ldots,p$. Then the matrix $B_{N-1}^{\prime}$ has eigenvalues $\eta_{i}$ with multiplicity $N_{i}$ for every $i\in\{0,\ldots,p\}\setminus\{k\}$, and $\eta_{k}$ with multiplicity $N_{k}-1$. Moreover, the smallest eigenvalue $\Lambda_{1}$ of $B_{N-1}^{\prime}$ equals $\Phi(\Lambda,Y(v_{1}))$.
@@
Let us prove (1). Let us first show the continuity of $\Phi(\Lambda,Y)$ on $Y\subset[\varepsilon,1]\times[\kappa,1]\times[0,1]^{p-2}\cap\Delta^{p+1}$ for some $\varepsilon>0$. We see from (2.15) that $\Lambda_{1}$ stays away from $\Lambda$ and $\eta_{1}$. In fact, because $y_{0}\geq\varepsilon$ and $\sum y_{i}\leq 1$
$\frac{\varepsilon}{\Lambda_{1}-\Lambda}\leq\frac{y_{0}}{\Lambda_{1}-\Lambda}=\sum\frac{y_{k}}{\eta_{k}-\Lambda_{1}}\leq\frac{1}{\eta_{1}-\Lambda_{1}}$
implies that $\Lambda_{1}-\Lambda\geq\varepsilon(\eta_{1}-\Lambda_{1})$ which implies $\Lambda_{1}-\Lambda\geq\varepsilon(1+\varepsilon)^{-1}(\eta_{1}-\Lambda)$. The same argument shows that since $y_{1}\geq\kappa$ and $y_{0}\leq 1$,
$\frac{1}{\Lambda_{1}-\Lambda}\geq\frac{y_{0}}{\Lambda_{1}-\Lambda}=\sum\frac{y_{k}}{\eta_{k}-\Lambda_{1}}\geq\frac{y_{1}}{\eta_{1}-\Lambda_{1}}\geq\frac{\kappa}{\eta_{1}-\Lambda_{1}}$
shows that $\eta_{1}-\Lambda_{1}\geq\kappa(1+\kappa)^{-1}(\eta_{1}-\Lambda)$. Hence, $\Lambda_{1}$ is at distance of order $\varepsilon$ of $\Lambda$ and $\kappa$ of $\eta_{1}$. Thus, if $Y,Y^{\prime}\in[\varepsilon,1]\times[\kappa,1]\times[0,1]^{p-2}\cap\Delta^{p+1}$, we easily deduce from (2.15) that $\Phi(\Lambda,Y)$ is the zero of a smooth function and hence that there exists a finite constant $C(\varepsilon,\kappa)$ so that
$|\Phi(\Lambda,Y)-\Phi(\Lambda,Y^{\prime})|\leq C(\varepsilon,\kappa)\sum_{i=1}^{p}|y_{i}-y_{i}^{\prime}|\,,$
which implies the continuity of $\Phi(\Lambda,.)$ on $[\varepsilon,1]\times[\kappa,1]\times[0,1]^{p-2}\cap\Delta^{p+1}$. When $y_{0}$ goes to zero, it is clear that $\Phi(\Lambda,Y)$ was extended by continuity to $\Lambda$, which in turn completes the proof of the continuity of $\Phi(\Lambda,.)$ on $[0,1]\times[\kappa,1]\times[0,1]^{p-1}\cap\Delta^{p+1}$.
We next turn to the proof of (2). Let us show that $\eta_{k}$ has multiplicity $N_{k}-1$. The eigenspace of $B_{N-1}^{\prime}$ associated to $\eta_{k}$ is
@@
Since $v_{1}\notin E_{\eta_{k}}$, $E_{\eta_{k}}^{\prime}$ has dimension $N_{k}-1$. Moreover $\Lambda$ is not an eigenvalue of $B_{N-1}^{\prime}$. There are therefore $p$ new eigenvalues $\Lambda_{1}\leq\cdots\leq\Lambda_{p}$. Recall that if $\lambda_{1}\leq\cdots\leq\lambda_{N}$ are the eigenvalues of $B_{N}$ and $\lambda_{1}^{\prime}\leq\cdots\leq\lambda_{N-1}^{\prime}$ are the eigenvalues of $B_{N-1}^{\prime}$, then by the Courant-Fischer theorem, $\lambda^{N}$ and $(\lambda^{\prime})^{N}$ are interlaced:
@@
Therefore for every $i=1,\ldots,p$, we have $\Lambda_{i}\in(\eta_{i-1},\eta_{i}).$ Moreover, by the Courant-Fischer theorem, $\Lambda_{1}$ can be expressed as
@@
Let $u\in\mathbb{R}^{N}$ be a minimizer in the above problem. By variational calculus, there exist $C_{1},C\in\mathbb{R}$ such that
$B_{N}u=Cu+C_{1}v_{1}.$
Since $\langle u,v_{1}\rangle=0$, we have
$\langle u,B_{N}u\rangle=C=\Lambda_{1}.$
Since $\Lambda_{1}\notin\{\lambda_{1},\ldots,\lambda_{N}\}$, we deduce that for every $i=1,\ldots,N$, if we denote by $w_{i}$ the eigenvector of $B_{N}$ with eigenvalue $\lambda_{i}$ and for a vector $r\in S_{N-1}$ $r_{i}=\langle r,w_{i}\rangle$,
$u_{i}=C_{1}\frac{(v_{1})_{i}}{\lambda_{i}-\Lambda_{1}}.$
Thus, $\Lambda_{1}$ satisfies
$\langle u,v_{1}\rangle=\sum_{i=1}^{N}\frac{(v_{1})_{i}^{2}}{\lambda_{i}-\Lambda_{1}}=0,$